Skip to main content

IODP Expeditions 367 and 368

South China Sea Rifted Margin


Daily Science Report for 27 April 2017

Location: Hole U1502A (proposed Site SCSII-17A, 18°27.8720′N, 116°13.8381′E)

Science Update: We continued to RCB core Hole U1502A. Cores U1502A-12R to 23R sampled from 471 to 586 m, recovering 31 m (27%). Recovery was highly variable and formation dependent. In fine-grained intervals (Cores 12R–14R, 471–499.8 m, and Core 18R, 528.6–538.2 m), we had higher recovery (76%) and steady, relatively slow rates of penetration (ROP ~14 m/hr). In coarser intervals (Cores 15R–17R, 499.8–528.6 m, and 19R–23R, 538.2–585.9 m), inferred to be loosely consolidated sands, we had very low recovery (3%) and very fast penetration rates (ROP ~110 m/hr).

Cores U1502A-9R to 14R have been split and described, and they predominantly consist of late Miocene greenish gray nannofossil-rich clay interbedded with dark grayish brown clay with nannofossils. Calcareous microfossils are absent (probably dissolved) from Cores U1502A-13R and 14R. Bioturbation varies from moderate to high. Scientists are collecting and measuring data from Hole U1502A and submitting the Site U1501 reports. We held a science summary meeting for Hole U1501D, with the participation of Expedition 367 scientists.


Daily Science Report for 26 April 2017

Location: Hole U1502A (proposed Site SCSII-17A, 18°27.8720′N, 116°13.8381′E)

Science Update: We started the day drilling ahead in Hole U1502A, penetrating from 250.4 to 375 m. We retrieved the center bit, dropped a core barrel, and started RCB coring at 0700 h, retrieving Cores U1502A-2R to 11R from 375 to 471 m (28.2 m recovery, 29.4%). Core recovery was very low (1%–9%) in Cores 3R–8R, which are inferred to be unconsolidated sand layers interbedded with relatively consolidated silty clay, based on similar results at Site U1499, which was drilled on the same ridge and at similar water depth during Expedition 367. Recovery markedly improved in Cores 9R–11R (73%–95%) when we encountered a substantial formation change to a moderately consolidated silty clay between 442.2 and 471 m.

Cores U1502A-2R to 8R were split, measured, and described today. They consist of late Miocene interbedded dark gray clay with silt, dark greenish gray silty clay, and clay-rich nannofossil ooze, with abundant trace fossils. Scientists started collecting data from Hole U1502A and submitting Site U1501 reports.


Daily Science Report for 25 April 2017

Location: Hole U1502A (proposed Site SCSII-17A, 18°27.8720′N, 116°13.8381′E)

Science Update: We started the day lowering the drill string to the seafloor. Before we reached bottom, we conducted required routine rig servicing (drill line slip and cut), and lowered the subsea camera system to observe the bit tag the seafloor to establish the water depth (3763.7 m). Hole U1502A was spudded at 1300 h. We spent the rest of the day drilling ahead with a center bit in place, and by midnight the drill bit had reached 250.4 mbsf. We plan to drill to 380 m, retrieve the center bit, and start RCB coring to determine a good interval in the formation to set the bottom of the casing. At our noon crossover, the Co-Chief Scientists presented the scientific objectives and operations plan for Site U1502. Scientists continued collecting, analyzing, and writing up the data from Site U1501.


Daily Science Report for 24 April 2017

Location:
Hole U1501D (proposed Site SCSII-41A, 18°53.0929′N, 115°45.9370′E)
In transit
Arrived at Site U1502 (proposed Site SCSII-17A, 18°27.8718′N, 116°13.8378′E, water depth 3761 mbsl)

Science Update: We started the day pulling out of Hole U1501D and positioning the end of the pipe at 113 m in preparation for downhole logging. The triple combo tool string was rigged up and lowered through the drill string and out into the open hole. As it descended into the hole, the bottom of the tool string encountered an obstruction at 299.3 mbsf. We raised and lowered the tool string was a few times in an attempt to pass through this obstruction, without success. We conducted our main pass upwards from 299.3 m, collecting magnetic susceptibility, resistivity, density (with caliper), neutron porosity, and gamma radiation data. At 156 mbsf, it was necessary to apply ~1000 lb of overpull to free the tool as the borehole diameter had been reduced to only 6 inch based on the caliper reading. Because of the obstruction and signs that the hole was caving in, we decided to cancel all remaining logging runs. The tool string was rigged down and the drill string was pulled up. After the rig floor was secured, we raised the thrusters, and started our transit to Site U1502 (proposed Site SCSII-17A) at 1800 h. At 2145 h, we arrived over the site coordinates. We lowered the thrusters, deployed the seafloor positioning beacon, and by midnight we had started making up the RCB bottom-hole assembly to start operations in Hole U1502A. Scientists continued collecting, analyzing, and writing up the data from Site U1501.


Daily Science Report for 23 April 2017

Location: Hole U1501D (proposed Site SCSII-41A, 18°53.0929′N, 115°45.9370′E)

Science Update: We completed coring operations in Hole U1501D with RCB Cores 24R to 27R sampling from 625.2 m to the final depth of 644.3 m at 1600 h. We recovered 14.7 m (77.1%) of poorly sorted, feldspar-rich, coarse to very coarse sandstone, interbedded with intervals of finely and coarsely laminated, well-sorted fine to medium-grained and strongly lithified sandstone. Following the retrieval of the last core we started preparing the hole for wireline logging. We circulated the hole with mud and conducted a series of wiper trips and mud sweeps to ensure the hole was clear. Afterward, we dropped the bit at the bottom of the hole and started to pull the end of the pipe to logging depth by midnight. Age determinations in Hole U1501D continued to be problematic because of a lack of planktonic microfossils. Scientists continue collecting, analyzing, and writing up the data from Site U1501. We held the second part of the science summary meeting for Holes U1501A–U1501C, with the participation of Expedition 367 scientists who joined the meeting using the shipboard videoconferencing system.


Daily Science Report for 22 April 2017

Location: Hole U1501D (proposed Site SCSII-41A, 18°53.0929′N, 115°45.9370′E)

Science Update: RCB coring in Hole U1501D penetrated from 548.6 to 625.2.1 m with an overall recovery of 42.71 m (Cores U1501D-14R to 23R; 55.8%). Core recovery continued to be highly variable due to the alternation of thin (meter-thick) sandstone layers and intervals inferred to be of slightly consolidated sand. It improved considerably after 606 m (Cores 20R to 23R; >80% recovery), when we encountered massive sandstone, and we decided to change the coring strategy to pulling the core barrel at half core length (~4.7 m). However, the harder formation also slowed the penetration rates to ~1 m/hr.

Cores U1501D-6R to 20R were split, measured, and described. Cores 6R to 18R are interbedded calcite-rich fine- to coarse-grained sandstone with abundant quartz and dark gray organic-rich and slightly lithified clayey silt with occasional lignite. Cores 19R to 20R sampled poorly sorted feldspar-rich coarse-grained sandstone, with a cobble-size rounded granite. An age determination for these cores has not been possible because they are barren of planktonic microfossils. Continuing micropaleontological examination of the basal sediment in Hole U1501C yielded a middle Eocene age for the base of Core 49X (350 mbsf). Scientists continue collecting, analyzing, and writing up the data from Site U1501, and preparing presentations for tomorrow’s summary meeting.


Daily Science Report for 21 April 2017

Location: Hole U1501D (proposed Site SCSII-41A, 18°53.0929′N, 115°45.9370′E)

Science Update: After drilling ahead without coring from 327.6 to 433.5 m, we retrieved the center bit, circulated mud to clean cuttings out of the hole, and started RCB coring at 0545 h. Cores U1501D-2R to 13R arrived aboard before midnight, sampling from 433.5 to 548.6 m, and recovering 21.33 m (18.5%). Recovery was low as we cored through thin, hard sandstone beds interbedded within thick, soft unconsolidated layers.

Cores U1501D-2R to 5R were split, measured, and described today. Cores 2R and 3R yielded alternating beds of clayey silt with sand and silty sand with abundant sand-size quartz and glauconite, and small shell fragments. Cores 4R and 5R revealed massive fine- to coarse-grained sandstone, with cobble-size clasts, abundant quartz, and other felsic minerals. Small mollusk shell fragments and benthic foraminifer tests are also present. Calcareous nannofossil biostratigraphy provides a preliminary age estimate of late Eocene for the base of Core 5R. Scientists continue collecting, analyzing, and writing up the data from Holes U1501A–U1501C. From 1130–1230 h, we held a meeting to present the first half of the results from Holes U1501A–U1501C. We hosted at least 10 scientists from the previous Expedition 367, who joined the meeting using the shipboard videoconferencing tools. A second meeting to present the remaining results will take place at the same time on Sunday 23 April.


Daily Science Report for 20 April 2017

Location: Hole U1501D (proposed Site SCSII-41A, 18°53.0928′N, 115°45.9369′E)

Science Update: The crew successfully completed replacement of the drawworks low clutch diaphragm at 0500 h. Following the repair, the crew reassembled the RCB BHA and the drill collars and lowered the drill string to the seafloor. Hole U1501D was spudded at 1305 h, 20 m west of Hole U1501C. We spent the rest of the day drilling ahead with a center bit in place, and by midnight the drill bit had reached 327.6 mbsf. We plan to drill to 433.5 m into the seafloor, retrieve the center bit, and then start RCB coring to our target depth.

The last cores from Hole U1501C (57X to 62X) were described today. Cores 57X to 59X (417.6 to 446.3 m) yielded alternating beds of glauconite within dark greenish gray to very dark greenish gray silty sand and clayey silt, with abundant benthic foraminifers and small mollusk shell fragments. Bioturbation is slight to moderate. Cores 60X to 62X (446.3 to 461.9 m) contain dark greenish gray sand and dark gray to black sandstone. Scientists continue to collect, analyze, and write up the data from Holes U1501A–U1501C, and to prepare presentations for tomorrow’s science summary meeting.


Daily Science Report for 19 April 2017

Location: Hole U1501C (proposed Site SCSII-41A, 18°53.0919′N, 115°45.9485′E)

Science Update: We completed XCB coring in Hole U1501C at 0750 h with Cores 59X to 62X sampling from 436.8 to 461.9 m, and recovering 19.49 m (78%). XCB core recovery diminished considerably in the last three cores, with Core 62X yielding only 46%. Thus, we made the decision to stop XCB coring and to end Hole U1501C. We pulled out of the hole and the APC/XCB bit was back on the rig floor at 1420 h. While making up the RCB BHA for rotary drilling/coring in Hole U1501D, the low clutch diaphragm on the drawworks failed at 1600 h. With the low clutch inoperable, it was not be possible to lift the drill string from depth due to heavy load. The crew secured all BHA components at the rigfloor and began repairs immediately. Repairs continued for the rest of the day.

Cores U1501C-46X to 56X were split, measured, and described today. Cores 46X–52X consist of dark greenish gray nannofossil-rich clay with sparse glauconite. The formation changes into very dark greenish glauconite-rich silty clay, clayey silt, and silty sand with abundant benthic foraminifera and mollusk shell fragments in Cores 53X–56X. Calcareous nannofossils provide an age estimate of early Oligocene for this sequence. Scientists continued analyzing samples, and they are starting to write the results from Holes U1501A–U1501C.


Daily Science Report for 18 April 2017

Location: Hole U1501C (proposed Site SCSII-41A, 18°53.0919′N, 115°45.9485′E)

Science Update: We continued to deepen Hole U1501C with Cores 43X to 58X sampling from 283.3 to 436.8 m and recovering 152 m (99%). XCB core recovery was very high for each XCB core, averaging 101% except for Core 53X, which only recovered 6.1 m (64%). Inspection of this core upon retrieval showed that a piece of sandstone was jammed into the coring shoe, preventing collection of any more material in the barrel.

Cores U1501C-20F to 45X were split, measured, and described today. They yielded predominantly heavily bioturbated, dark greenish gray to light brownish gray, clay-rich nannofossil ooze and nannofossil ooze with foraminifers. An early Oligocene age has been assigned to the base of Core 54X. Scientists continued to collect and analyze core samples, and they were given a presentation about depth scales during our noon general crossover meeting.


Daily Science Report for 17 April 2017

Location: Hole U1501C (proposed Site SCSII-41A, 18°53.0919′N, 115°45.9485′E)

Science Update: We started the day by retrieving Core U1501C-17H from 156.8 mbsf. Because this core was only a partial stroke, we advanced the bit by recovery (4.95 m or 99%), then switched to the half-length APC system (HLAPC). We continued deepening the hole, and Cores U1501C-18F to 39F sampled from 156.8 m to HLAPC refusal at 257 m (88.7 m recovery; 88.6%). Then we switched to the XCB coring system. Cores U1501C-40X to 42X extended the hole from 257 to 283.3 m, recovering 28 m of core (106%).

Cores U1501C-4H to 19F were split, measured, and described today. They consist of dark greenish gray to light brownish gray nannofossil ooze with foraminifera and clay-rich nannofossil ooze. Bioturbation varies from moderate to heavy. An early Miocene age has been assigned to the base of Core 33F.


Daily Science Report for 16 April 2017

Location:
Hole U1501A (proposed Site SCSII-41A, 18°53.0921′N, 115°45.9426′E)
Hole U1501B (proposed Site SCSII-41A, 18°53.0921′N, 115°45.9426′E)
Hole U1501C (proposed Site SCSII-41A, 18°53.0919′N, 115°45.9485′E)

Science Update: We started the day by retrieving Core U1501A-1H from 2868 mbrf. The mud line core recovered 9.73 m of sediment, and it was determined that the piston core had fired from below the seafloor. Hole U1501A was terminated at 0025 h. The drill string was repositioned to attempt to core a better mud line at the site. With the bit at 2663 mbrf, we started Hole U1501B at 0100 h, again recovering a full core barrel, indicating that the APC was fired at or below the seafloor, and Hole U1501B was terminated. We offset the vessel 10 m to the east and raised the bit to 2843 mbrf to start coring in Hole U1501C. This and the following attempt from 2848 mbrf retrieved empty core barrels. On a third attempt to start coring, the bit was lowered to 2856.4 mbrf, and Hole U1501C was started at 0510 h. Core U1501C-1H recovered 9.2 m of sediment (99%) and established the seafloor at 2845.8 mbsl. APC Cores 1H–17H penetrated from the seafloor to 156.8 m and recovered 156.6 m (100%). All APC cores were oriented, and we made formation temperature measurements (APCT-3) on Cores U1501C-5H (47.3 m), 8H (75.8 m), 11H (94.8 m), and 14H (132.8 m). Cores U1501A-1H, U1501B-1H, and U1501C-1H to 3H have been split and described; they are predominantly dark greenish gray clay-rich nannofossil ooze with foraminifera. Bioturbation ranges from moderate to heavy. Biostratigraphic datums indicate a late Miocene age for the base of Core U1501C-16H.


Daily Science Report for 15 April 2017

Location:
In transit
Site U1501 (proposed Site SCSII-41A, 18°53.0922′N, 115°45.9426′E; water depth 2846 mbsl)

Science Update: After a 255 nmi transit from Hong Kong we arrived at Site U1501 (Prospectus Site SCSII-41A) at 0800 h. We lowered the thrusters, deployed a positioning beacon 50 m west of the site coordinates, put together the APC/XCB bottom-hole assembly, and started lowering it to the seafloor in preparation for coring. Pipe tripping continued throughout the day until 2200 h when we installed the sinker bars and the core orientation tool, and adjusted the drill string length to position the bit for taking the first core. Coring in Hole U1501A started at 2345 h. During the day, the Expedition Project Manager/Staff Scientist and the Assistant Laboratory Officer introduced the scientists to the shipboard applications used to retrieve and view core data. Scientists continue to learn the laboratories using thin sections, smear slides, and residues from Expedition 367.


Daily Science Report for 14 April 2017

Location:
China Merchants Wharf, Hong Kong
In transit to Site U1501 (proposed site SCSII-41A)

Science Update: Throughout the early morning, the ship’s crew completed securing all equipment. After immigration activities were completed, the last line was released at 0912 h, ending all portcall activities and beginning our sea passage. The Co-Chief Scientists gave a presentation on the Site U1501 science objectives and what we expect to encounter. In the afternoon, the Captain held the first fire and boat safety drill. The scientists resumed laboratory team meetings and moved on to their working shifts. The rest of the day was spent in transit to Site U1501 (proposed Site SCSII-41A).


Daily Science Report for 13 April 2017

Location: China Merchants Wharf, Hong Kong

Science Update: The day started with the South China Sea Expedition 368 scientists explaining their individual research objectives. Later, the Co-Chief Scientists, the Staff Scientist, and the Shipboard Curator met to discuss and approve the shipboard sampling plan. The laboratory groups continued learning their laboratories and preparing their instruments and methods. The scientists were given the afternoon off because our departure is scheduled for tomorrow, 14 April. We completed loading 131 joints of 10¾ inch casing and two hydraulic release tool (HRT) reentry systems, and we offloaded 204 joints of 5½ inch drill pipe for inspection.


Daily Science Report for 12 April 2017

Location: China Merchants Wharf, Hong Kong

Science Update: In the morning, scientists were introduced to drilling and coring operations by the JRSO Operations Superintendent, and to the downhole logging tools by the Logging Engineer. Our Educators and Journalists gave an overview of their education and outreach plans for the expedition. In the afternoon, the Shipboard Curator and Expedition Project Manager introduced the scientists to core flow and sampling, and the laboratory groups continued learning their laboratories and preparing their instruments.


Daily Science Report for 11 April 2017

Location: China Merchants Wharf, Hong Kong

Science Update: The day began with introductions of the Expedition 368 scientists and JRSO technical staff, followed by a presentation of the scientific objectives by the Co-Chief Scientists and a presentation of the shipboard deliverables by the Expedition Project Manager. In the afternoon, after the Captain’s introduction and safety orientation, scientists broke up into laboratory groups to meet their JRSO technical staff team members.

A faulty detector in the natural gamma ray (NGR) system was replaced, and diagnostics indicate all systems are functional. The superconducting rock magnetometer (SRM) calibration was completed successfully by the vendor. The technician is working with the JRSO technical staff to improve diagnostics. The sample preparation system for inductively coupled plasma spectroscopy (ICP) analyses (i.e., bead maker) was repaired, and it is functional.


Daily Science Report for 10 April 2017

Location: China Merchants Wharf, Hong Kong

Science Update: The Expedition 368 scientists boarded the ship in the morning and got settled in their rooms. Following this, they were introduced to life aboard the JOIDES Resolution, received a general safety presentation, and participated in an initial laboratory and ship safety tour. In the afternoon, the scientists were introduced to the shipboard computing environment and connected their computers to the ship’s network. Transfer of incoming and outgoing shipments continued throughout the day.


Daily Science Report for 9 April 2017

Location: China Merchants Wharf, Hong Kong

Science Update: The South China Sea Rifted Margin Expedition 368 started at 0754 h with the first line ashore at the China Merchants Wharf in Hong Kong. The Co-Chief Scientists, a crossover science team from Expedition 368, and IODP staff moved onto the ship and started crossover with their Expedition 367 counterparts. Initial loading of incoming shipments began.